OpenStudy (anonymous):

Two times the difference between a number and eight is less than or equal to five times the sum of the number and four. Find the smaales number that will satisfy the inequality.

OpenStudy (anonymous):

2(n-8)<=5(n+4) 2n-16<=5n+20 -3n<=36 n>=-12

OpenStudy (anonymous):

so n=-12

OpenStudy (anonymous):

Thanks myko :-)

OpenStudy (anonymous):

yw